OpenStudy (anonymous):

Pls help Use the unit rate to determine how many cups of nuts are needed to make 1 lb of fudge. Show your work. 9/4 cups of nuts to make 1/2 lbs of fudge 9/4 / 1/2 (complex fraction) How do I change it into a unit rate

OpenStudy (anonymous):

\[\frac{ 9/4 }{ 1/2 }\]

OpenStudy (anonymous):

\[\frac{ 9 }{ 4 } \over \frac{ 1 }{ 2}\]

OpenStudy (anonymous):

complex fraction, pls help solve for unit rate

OpenStudy (mathstudent55):

Do you know how to divide fractions?

OpenStudy (anonymous):

Do I have to change the fractions into decimals?

OpenStudy (mathstudent55):

\(\Large \dfrac{\frac{9}{4}}{\frac{1}{2}} = \dfrac{9}{4} \div \dfrac{1}{2} \)

OpenStudy (anonymous):

oh, nvm lol

OpenStudy (anonymous):

9/2?

OpenStudy (mathstudent55):

You don't need decimals. Your complex fraction is a fraction divided by a fraction. Just treat it that way. Multiply the first fraction by the reciprocal of the second fraction.

OpenStudy (anonymous):

9 over 2 ?

OpenStudy (anonymous):

@mathstudent55 I need help...?

OpenStudy (anonymous):

Thanks for the help I got 4 1/2 over 1 :I

OpenStudy (kayders1997):

When you divide fractions you can switch the second fraction and multiply

OpenStudy (anonymous):

Ik

OpenStudy (anonymous):

congratulations

OpenStudy (mathstudent55):

Yes, the unit rate is is 9/2 cups of nuts per lb of fudge. \(\Large \dfrac{ \frac{9}{4} }{ \frac{1}{2} } = \dfrac{9}{4} \div \dfrac{1}{2} = \dfrac{9}{4} \times \dfrac{2}{1} = \dfrac{18}{4} = \dfrac{9}{2} \)